1. For the given random process of the total number of visitors that visit a museum per hour, find the followings: 1415 p(x)0.150.120.110.20.080.090.080.050.12 2 10 12 a) Mean, variance, and standard deviation. b) The probability that the total number of visitors not exceed 3 visitors. c) The probability that the total number of visitors not less than 9 visitors
Answer #1

(a) mean=E(x)=sum(x*p(x))=7.2 ,

variance=E(x2)-E(x)*E(x)=76.98-7.2*7.2=25.14

standard deviation=sqrt(vaiance)=sqrt(25.14)=5.014

x p(x) x*p(x) x2*p(x)
1 0.15 0.15 0.15
2 0.12 0.24 0.48
5 0.11 0.55 2.75
7 0.2 1.4 9.8
9 0.08 0.72 6.48
10 0.09 0.9 9
12 0.08 0.96 11.52
14 0.05 0.7 9.8
15 0.12 1.8 27
sum= 1 7.42 76.98

(b) P(not exceed 3)=P(X=1)+P(x=2)=0.15+0.12=0.27

(c)P(not less than 9)=P(X=10)+P(X=12)+P(X=14)+P(X=15)=0.09+0.08+0.05+0.12=0.34